# See pages 1 of figures/front.pdf

### v1

@@ Introduction
In this chapter we look at high-energy particle physics at the LHC, specifically through the lens of the LHCb experiment, with which this work is associated. The detector of the experiment, the dataflow and its trigger system are described. Only the Upgrade I detector for Run 3 [216] of the LHC is discussed. Information about the previous LHCb configuration can be found in [217, 218].
@@ 6.1 The Large Hadron Collider at CERN
The LHC [220] at CERN is currently the world’s biggest and most powerful particle accelerator located roughly 100 meters below ground. The accelerator has the form of a ring with a perimeter of 27 km and 90% of its length is in molasse rock while the rest 10% is in limestone under the Jura mountain. This machine is mainly used to accelerate protons as well as heavy ions, such as lead, in order to collide them. The protons are accelerated in bunches, as illustrated in Fig. 6.2, in two superconducting magnet rings with opposite directions. After the particles have been accelerated, they are brought into collision at four interaction points hosting the detectors for the four main CERN experiments: ATLAS, CMS, ALICE and LHCb, featured on Fig. 6.3. These bunches of particles cross every 25 ns, or equivalently at a frequency of 40 MHz.
@@
LHC operates in periods called “Runs”. Run 1 took place between 2010–2012, Run 2 between 2015–2018 and Run 3 started in July 2022. Run 3 is planned to last until July of 2026 while Run 4 is scheduled for the summer of 2030. The LHC is designed to accelerate protons very close to the speed of light, reaching energies up to 7 TeV. During Run 3, the LHC collides protons with a center-of-mass energy of $\sqrt{s}=$ 13.6 TeV [^1] Variable $s$ is the Mandelstam $s$ variable [223] defined as $s=(\mathbf{p}_{1}+\mathbf{p}_{2})^{2}c^{2}$, where $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ are the four-momenta of the incoming particles. $\sqrt{s}$ is the center-of-mass energy, and is an observer-independent way to measure how hard the protons are being collided against each other.. While ATLAS and CMS operate at a peak instantaneous luminosity of $\mathcal{L}=$ $2\times 10^{34}$ cm-2s-1 which decreases with time during an LHC fill, LHCb uses luminosity leveling [224] in order to keep the luminosity lower but constant. This is done in order to deliver steady conditions for physics analysis, but also because the processes that LHCb focuses on are difficult to record at high luminosities. For Run 3, LHCb aims for a instantaneous luminosity of $\mathcal{L}=2\times 10^{33}$ cm-2s-1, which results in about five pppp collisions per bunch crossing on average.
@@ 6.3 Vertex Locator
The VELO modules are being cooled using evaporative bi-phase CO2 [237] absorbing the heat generated by the VeloPix read-out chips in order to ensure stable performance. The sensors operate in a secondary vacuum, which is separated from the LHC beam vacuum by an aluminum Radio Frequency (RF) box [238]. The RF box further shields the detector electronics from RF pickup of the beams. The RF foil needs to be mechanically stable, in order to withstand pressure changes and differences, and extremely radiation hard due to the intense radiation environment very close to the LHC beam.
@@ 6.4 Online System and Data Acquisition
The LHCb Data Acquisition (DAQ) system, illustrated in Fig. 6.9, consists of 162 Event Builder (EB) servers with AMD EPYC 7502 32-core CPUs hosting custom-made FPGA detector read-out boards (TELL40) and Nvidia RTX A5000 GPUs on the Peripheral Component Interconnect Express (PCIe) interface. The DAQ design is “triggerless”, meaning that the detector is read out at the nominal LHC bunch-crossing frequency of 40 MHz. The more recently updated layout, with three GPU cards running HLT1 on the EB servers, and with the EB output data rate reduced by a factor between 30 and 60, is shown in Fig. 6.10.
@@ 6.7 LHCb Trigger System
As already discussed in Chapter 2, the amount of data produced by each of the main four LHC experiments is too massive to be stored on disk or tape. For example, a typical LHCb event under Run 3 conditions is around 100 kB of data, which at the LHC $pp$ collision rate of 40 MHz results in a data rate of 4 TB/s. In 2017, for Run 2, the LHC delivered stable beams for 1634 hours [257], so assumming the same availability for Run 3, the amount of data accumulates to 18 EB ($10^{18}$ B) per year. Together with the other three LHC experiments, the sum ends up at around $\mathcal{O}(100)$ EB per year, a number comparable to the monthly global internet traffic [258] and unfeasible to store permanently. This is the reason why all large HEP experiments have sophisticated triggers systems that filter the interesting collision events.
@@
Figure 6.16: Illustration of the LHCb trigger computing model. The same reconstructed event is saved with varying levels of object persistence: turbo (top), selective persistence (middle), and complete recostruction persistence (bottom). Top: A candidate $D^{0}\to K^{-}\pi^{+}$ is selected by HLT2; only the candidate and the Primary Vertex (PV) are persisted. Middle: Additional objects, e.g., pion tracks from candidate $D^{*+}\to D^{0}\pi^{+}$, can also be persisted. Bottom: The full reconstruction event is persisted, including raw subdetector data banks. Solid lines and objects denote persisted information in each case. Raw banks are represented as rectangles. Figure from [266, 262].
@@ 8.1 Early Version of ETX4VELO
In our problem, the hits left by charged particles in the VELO detector form a point cloud in 3-dimensional space. However, they can also be represented as a graph, in which the successive hits (or nodes) of each particle traversing the detector are connected to each other. This graph can be thought of as the “truth graph” that perfectly describes the VELO event. At a high level, the end-goal of the ETX4VELO pipeline is to produce a graph that closely approximates this truth graph.
@@
At this point the architecture of the embedding MLP was 4 hidden layers of 256 neurons, ReLU [312] activations, and with about 200 000 total number of parameters. On the other had, the GNN had roughly two million parameters.
An example of the evaluation of the pipeline on 5000 events using MonteTracko is shown in Fig. 8.4. MonteTracko is the custom evaluation suite developed specifically with the ETX4VELO pipeline in mind. It offers tools for matching simulated particle trajectories with reconstructed tracks and calculating performance metrics. TrackChecker, on the other hand, is the native set of algorithms used to check the tracks created by the Allen tracking system against the Monte Carlo (MC) truth information from the simulation. Similarly to the TrackChecker evaluation, the MonteTracko output is split across various particle categories—01_velo, 02_long etc.—and across various track-finding performance metrics: clone rate (clones), ghost rate (ghosts), hit purity (pur) and hit efficiency (hit eff). Therefore, the output is identical to Allen’s VELO validation output.
@@ 8.2 Reconstruction of Electrons
The first naive attempt to solve this problem was to modify the training sample. For the rest of this section when we refer to electrons we mean both electrons and positrons, unless stated otherwise. One approach was to create events that contain only electrons, around 1500 each, by manually deleting all other particles, and training on them. The second approach was to select events that have a higher number of electrons, and then train on them. With these approaches the score for long electrons was improved to above 80%, without any modifications to the architecture. However, the problem, as it turned out later, is more fundamental and is not really related with the balancing of the dataset.
@@ 9.1 Development
The throughput, as expected, was quite low. Only the embedding step, using the architecture at the time, comprising 4 hidden layers of 256 neurons, achieved a throughput of 8000 events per second.
@@
The second approach was to use the Exa.TrkX implementation of their pipeline on C++ [307], which uses ONNX Runtime for the inference of the ML models, and integrate it into Allen. The focus now was to have something that works end to end, in order to be able to work on incrementally improving it. In this implementation ONNX Runtime is used to run the models on CPU, even though the initial design goal of the Exa.TrkX implementation was targeting GPU. By working on this implementation, I was able to further understand the intricacies of the C++ language, and this experience proved essential, later on, for implementing the various algorithms on GPU in CUDA.
@@
This end to end implementation, with the exported models from the early version of the pipeline, results in a throughput of around 20 events per second. Having started from the original Exa.TrkX pipeline, the models were minimally modified, and hence were extremely heavy at this point, resulting in low throughput. Particles within the ATLAS and CMS detectors exhibit curved and helical trajectories which means that the reconstruction models need to be significantly deeper in order to learn the corresponding patterns. In the VELO detector, however, due to the lack of any magnetic field, the particles move in straight lines and hence the pattern recognition task becomes simpler. Because of this, we were able to reduce the size of the models. The embedding MLP was reduced from around 200 000 to 251 parameters, while the GNN was reduced from two million parameters down to only 70 000, as shown in Fig. 9.2. With these smaller models, and after a series of optimizations of the algorithms for throughput, we arrive at the final ETX4VELO pipeline, which is presented next.
@@ 9.2.2 Network Inference
Inference for the embedding network and the GNN, both trained in PyTorch, is performed using inference engines. We experimented with ONNX Runtime (ORT) leveraging its CUDA backend, and TensorRT (TRT), as shown in Fig. 9.5. Both engines require the PyTorch models to be exported with ONNX, using the torch.onnx method. TensorRT integrates seamlessly with the Allen framework for memory management, making it particularly well-suited for real-time inference and production workflows. ONNX Runtime, on the other hand, does not natively support this feature but provides the flexibility of a CPU backend, enabling easy transitions between GPU and CPU-based pipelines. To set up the environment dependencies for model inference in Allen, a CUDA-enabled cross-compilation toolchain was used [343].
@@ 10.1.1 PYNQ Framework
Similarly to software libraries, which can be called by the programmer to perform certain tasks while avoiding the intricate details of the implementation, overlays can be loaded to the FPGA dynamically. For example, for an image processing application, the various image processing functions, such as edge detection, compression etc., could be implemented in different overlays and loaded from Python, as required.
@@ 10.1.3 Workflow
In HLS4ML, there are the concepts of the frontend and the backend. The frontend is responsible for parsing the input neural network into an internal model graph, while the backend determines the type of output generated from this graph. These frontends and backends can be selected independently. For example, frontends include parsers for Keras or ONNX, while backends include Vivado HLS, Intel HLS, and Vitis HLS. Here I chose the VivadoAccelerator backend. The VivadoAccelerator backend of HLS4ML leverages the PYNQ software stack to easily deploy models on supported devices. For this backend, the I/O type, the hardware part that is being targeted, the clock period etc. has to be specified. The target FPGAs for my implementations are the PYNQ-Z2 board, which contains a Xilinx Zynq-7020 FPGA, and the Alveo U50 and U250 featuring the UltraScale+ and XCU250 FPGAs, respectively. It is important to note that the PYNQ-Z2 board is designed for educational purposes, whereas the Alveo cards are significantly larger and intended for use in data centers. All three cards are supported by the PYNQ project.
@@ Conclusion
Finally, the pipeline was partially implemented on FPGAs. Since the GNN comprises multiple MLPs, the implementation focused one of its fundamental components: the MLP. Furthermore, due to time constraints and the scope of the challenge, the implementation of the GNN was deferred to future work. The embedding MLP, on 8 bits, was benchmarked against the GPU implementation, and the implementations were compared based on energy efficiency considerations.
The ETX4VELO work would be interesting to be extended to the other tracking detectors of the LHCb experiment: the SciFi or the UT. However, compared to the VELO, these detectors, and hence their tracking algorithms [437, 328, 438], have different configurations and operating principles, making such an extension non-trivial. Secondly, the triplet-based methodology, currently left out of the GPU implementation, could be implemented, and the impact on the computational and physics performance re-evaluated. Thirdly, once the compatibility and support issues have been resolved, the quantization of the GNN, could be pursued. This avenue provides the most hope for throughput gains for the ETX4VELO pipeline. Finally, the k-NN and WCC algorithms can potentially be redeveloped to leverage more efficiently the GPU resources.
@@
Furthermore, a complete implementation, instead of an extrapolation, of packing multiple embedding neural networks on Alveo cards, can be done. In this way, the large-scale, high-throughput environment needed in the context of an LHC trigger can be simulated, and the comparison between the two hardware architectures can be conducted more fairly, taking into account data transfers and I/O overheads.
The work related to the ETX4VELO pipeline can be further extended to other applications, such as primary vertex finding, building upon the work in [439, 440] or graph clustering for electromagnetic calorimeters [441]. The end of the pipeline can be minimally modified to match the task at hand, leaving the graph architecture unchanged, and trained accordingly to leverage the new datasets.
@@ Appendix B Early ETX4VELO Development
Using XDIGI2CSV, the VELO data were split into two files. The hits_velo.csv file contains the VELO cluster coordinates as well as the Monte Carlo (MC) identifier of the origin particle of this cluster. The file mc_particles.csv associates, for each event, the MC identifier with the properties of the corresponding particle, e.g., momentum, pseudorapidity, etc.

### v2

@@ Introduction
In this chapter, we look at high-energy particle physics at the LHC, specifically through the lens of the LHCb experiment, with which this work is associated. The detector of the experiment, the dataflow and its trigger system are described. Only the Upgrade I detector for Run 3 [216] of the LHC is discussed. Information about the previous LHCb configuration can be found in [217, 218].
@@ 6.1 The Large Hadron Collider at CERN
The LHC [220] at CERN is currently the world’s biggest and most powerful particle accelerator located roughly 100 meters below ground. The accelerator has the form of a ring with a perimeter of 27 km and 90% of its length is in molasse rock while the remaining 10% is in limestone under the Jura mountain. This machine is mainly used to accelerate protons as well as heavy ions, such as lead, in order to collide them. The protons are accelerated in bunches, as illustrated in Fig. 6.2, in two superconducting magnet rings with opposite directions. After the particles have been accelerated, they are brought into collision at four interaction points hosting the detectors for the four main CERN experiments: ATLAS, CMS, ALICE and LHCb, featured on Fig. 6.3. These bunches of particles cross every 25 ns, or equivalently at a frequency of 40 MHz.
@@
LHC operates in periods called “Runs”. Run 1 took place between 2010–2012, Run 2 between 2015–2018 and Run 3 started in July 2022. Run 3 is planned to last until July of 2026 while Run 4 is scheduled for the summer of 2030. The LHC is designed to accelerate protons very close to the speed of light, reaching energies up to 7 TeV. During Run 3, the LHC collides protons with a center-of-mass energy of $\sqrt{s}=$ 13.6 TeV [^1] Variable $s$ is the Mandelstam $s$ variable [223] defined as $s=(\mathbf{p}_{1}+\mathbf{p}_{2})^{2}c^{2}$, where $\mathbf{p}_{1}$ and $\mathbf{p}_{2}$ are the four-momenta of the incoming particles. $\sqrt{s}$ is the center-of-mass energy, and is an observer-independent way to measure how hard the protons are being collided against each other.. While ATLAS and CMS operate at a peak instantaneous luminosity of $\mathcal{L}=$ $2\times 10^{34}$ cm-2s-1 which decreases with time during an LHC fill, LHCb uses luminosity leveling [224] in order to keep the luminosity lower but constant. This is done in order to deliver steady conditions for physics analysis, but also because the processes that LHCb focuses on are difficult to record at high luminosities. For Run 3, LHCb aims for an instantaneous luminosity of $\mathcal{L}=2\times 10^{33}$ cm-2s-1, which results in about five pppp collisions per bunch crossing on average.
@@ 6.3 Vertex Locator
The VELO modules are being cooled using evaporative bi-phase CO2 [237] absorbing the heat generated by the VeloPix readout chips in order to ensure stable performance. The sensors operate in a secondary vacuum, which is separated from the LHC beam vacuum by an aluminum Radio Frequency (RF) box [238]. The RF box further shields the detector electronics from RF pickup of the beams. The RF foil needs to be mechanically stable, in order to withstand pressure changes, and extremely radiation-hard due to the intense radiation environment very close to the LHC beam.
@@ 6.4 Online System and Data Acquisition
The LHCb Data Acquisition (DAQ) system, illustrated in Fig. 6.9, consists of 162 Event Builder (EB) servers with AMD EPYC 7502 32-core CPUs hosting custom-made FPGA detector readout boards (TELL40) and Nvidia RTX A5000 GPUs on the Peripheral Component Interconnect Express (PCIe) interface. The DAQ design is “triggerless”, meaning that the detector is read out at the nominal LHC bunch-crossing frequency of 40 MHz. The more recently updated layout, with three GPU cards running HLT1 on the EB servers, and with the EB output data rate reduced by a factor between 30 and 60, is shown in Fig. 6.10.
@@ 6.7 LHCb Trigger System
As already discussed in Chapter 2, the amount of data produced by each of the main four LHC experiments is too massive to be stored on disk or tape. For example, a typical LHCb event under Run 3 conditions is around 100 kB of data, which at the LHC $pp$ collision rate of 40 MHz results in a data rate of 4 TB/s. In 2017, for Run 2, the LHC delivered stable beams for 1634 hours [257], so assuming the same availability for Run 3, the amount of data accumulates to 18 EB ($10^{18}$ B) per year. Together with the other three LHC experiments, the sum ends up at around $\mathcal{O}(100)$ EB per year, a number comparable to the monthly global internet traffic [258] and unfeasible to store permanently. This is the reason why all large HEP experiments have sophisticated triggers systems that filter the interesting collision events.
@@
Figure 6.16: Illustration of the LHCb trigger computing model. The same reconstructed event is saved with varying levels of object persistence: turbo (top), selective persistence (middle), and complete reconstruction persistence (bottom). Top: A candidate $D^{0}\to K^{-}\pi^{+}$ is selected by HLT2; only the candidate and the Primary Vertex (PV) are persisted. Middle: Additional objects, e.g., pion tracks from candidate $D^{*+}\to D^{0}\pi^{+}$, can also be persisted. Bottom: The full reconstruction event is persisted, including raw subdetector data banks. Solid lines and objects denote persisted information in each case. Raw banks are represented as rectangles. Figure from [266, 262].
@@ 8.1 Early Version of ETX4VELO
In our problem, the hits left by charged particles in the VELO detector form a point cloud in 3-dimensional space. However, they can also be represented as a graph, in which the successive hits (or nodes) of each particle traversing the detector are connected to each other. This graph can be thought of as the “truth graph” that perfectly describes the VELO event. At a high level, the end goal of the ETX4VELO pipeline is to produce a graph that closely approximates this truth graph.
@@
At this point the architecture of the embedding MLP was four hidden layers of 256 neurons, ReLU [312] activations, and with about 200 000 total number of parameters. On the other hand, the GNN had roughly two million parameters.
An example of the evaluation of the pipeline on 5000 events using MonteTracko is shown in Fig. 8.4. MonteTracko is the custom evaluation suite developed specifically with the ETX4VELO pipeline in mind. It offers tools for matching simulated particle trajectories with reconstructed tracks and calculating performance metrics. TrackChecker, on the other hand, is the native set of algorithms used to check the tracks created by the Allen tracking system against the Monte Carlo (MC) truth information from the simulation. Similarly to the TrackChecker evaluation, the MonteTracko output is split across various particle categories—01_velo, 02_long, etc.—and across various track-finding performance metrics: clone rate (clones), ghost rate (ghosts), hit purity (pur) and hit efficiency (hit eff). Therefore, the output is identical to Allen’s VELO validation output.
@@ 8.2 Reconstruction of Electrons
The first naive attempt to solve this problem was to modify the training sample. For the rest of this section when we refer to electrons we mean both electrons and positrons, unless stated otherwise. One approach was to create events that contain only electrons, around 1500 each, by manually deleting all other particles, and training on them. The second approach was to select events that have a higher number of electrons, and then train on them. With these approaches the score for long electrons was improved to above 80%, without any modifications to the architecture. However, the problem, as it turned out later, is more fundamental and is not really related to the balancing of the dataset.
@@ 9.1 Development
The throughput, as expected, was quite low. Only the embedding step, using the architecture at the time, comprising four hidden layers of 256 neurons, achieved a throughput of 8000 events per second.
@@
The second approach was to use the Exa.TrkX implementation of their pipeline on C++ [307], which uses ONNX Runtime for the inference of the ML models, and integrate it into Allen. The focus now was to have something that works end-to-end, in order to be able to work on incrementally improving it. In this implementation ONNX Runtime is used to run the models on CPU, even though the initial design goal of the Exa.TrkX implementation was targeting GPU. By working on this implementation, I was able to further understand the intricacies of the C++ language, and this experience proved essential, later on, for implementing the various algorithms on GPU in CUDA.
@@
This end-to-end implementation, with the exported models from the early version of the pipeline, results in a throughput of around 20 events per second. Having started from the original Exa.TrkX pipeline, the models were minimally modified, and hence were extremely heavy at this point, resulting in low throughput. Particles within the ATLAS and CMS detectors exhibit curved and helical trajectories, which means that the reconstruction models need to be significantly deeper in order to learn the corresponding patterns. In the VELO detector, however, due to the lack of any magnetic field, the particles move in straight lines and hence the pattern recognition task becomes simpler. Because of this, we were able to reduce the size of the models. The embedding MLP was reduced from around 200 000 to 251 parameters, while the GNN was reduced from two million parameters down to only 70 000, as shown in Fig. 9.2. With these smaller models, and after a series of optimizations of the algorithms for throughput, we arrive at the final ETX4VELO pipeline, which is presented next.
@@ 9.2.2 Network Inference
Inference for the embedding network and the GNN, both trained in PyTorch, is performed using inference engines. We experimented with ONNX Runtime (ORT) leveraging its CUDA backend and TensorRT (TRT), as shown in Fig. 9.5. Both engines require the PyTorch models to be exported with ONNX, using the torch.onnx method. TensorRT integrates seamlessly with the Allen framework for memory management, making it particularly well-suited for real-time inference and production workflows. ONNX Runtime, on the other hand, does not natively support this feature but provides the flexibility of a CPU backend, enabling easy transitions between GPU and CPU-based pipelines. To set up the environment dependencies for model inference in Allen, a CUDA-enabled cross-compilation toolchain was used [343].
@@ 10.1.1 PYNQ Framework
Similarly to software libraries, which can be called by the programmer to perform certain tasks while avoiding the intricate details of the implementation, overlays can be loaded to the FPGA dynamically. For example, for an image processing application, the various image processing functions, such as edge detection, compression, etc., could be implemented in different overlays and loaded from Python, as required.
@@ 10.1.3 Workflow
In HLS4ML, there are the concepts of the frontend and the backend. The frontend is responsible for parsing the input neural network into an internal model graph, while the backend determines the type of output generated from this graph. These frontends and backends can be selected independently. For example, frontends include parsers for Keras or ONNX, while backends include Vivado HLS, Intel HLS, and Vitis HLS. Here I chose the VivadoAccelerator backend. The VivadoAccelerator backend of HLS4ML leverages the PYNQ software stack to easily deploy models on supported devices. For this backend, the I/O type, the hardware part that is being targeted, the clock period, etc. has to be specified. The target FPGAs for my implementations are the PYNQ-Z2 board, which contains a Xilinx Zynq-7020 FPGA, and the Alveo U50 and U250 featuring the UltraScale+ and XCU250 FPGAs, respectively. It is important to note that the PYNQ-Z2 board is designed for educational purposes, whereas the Alveo cards are significantly larger and intended for use in data centers. All three cards are supported by the PYNQ project.
@@ Conclusion
Finally, the pipeline was partially implemented on FPGAs. Since the GNN comprises multiple MLPs, the implementation focused on one of its fundamental components: the MLP. Furthermore, due to time constraints and the scope of the challenge, the implementation of the GNN was deferred for future work. The embedding MLP, on 8 bits, was benchmarked against the GPU implementation, and the implementations were compared based on energy efficiency considerations.
The ETX4VELO work would be interesting to be extended to the other tracking detectors of the LHCb experiment: the SciFi or the UT. However, compared to the VELO, these detectors, and hence their tracking algorithms [437, 328, 438], have different configurations and operating principles, making such an extension non-trivial. Secondly, the triplet-based methodology, currently left out of the GPU implementation, could be implemented, and the impact on the computational and physics performance re-evaluated. Thirdly, once the compatibility and support issues have been resolved, the quantization of the GNN could be pursued. This avenue provides the most hope for throughput gains for the ETX4VELO pipeline. Finally, the k-NN and WCC algorithms can potentially be redeveloped to leverage more efficiently the GPU resources.
@@
Furthermore, a complete implementation, instead of an extrapolation of packing multiple embedding neural networks on Alveo cards, can be done. In this way, the large-scale, high-throughput environment needed in the context of an LHC trigger can be simulated, and the comparison between the two hardware architectures can be conducted more fairly, taking into account data transfers and I/O overheads.
The work related to the ETX4VELO pipeline can be further extended to other applications, such as primary vertex finding, building upon the work in [439, 440], or graph clustering for electromagnetic calorimeters [441]. The end of the pipeline can be minimally modified to match the task at hand, leaving the graph architecture unchanged, and trained accordingly to leverage the new datasets.
@@ Appendix B Early ETX4VELO Development
Using XDIGI2CSV, the VELO data were split into two files. The hits_velo.csv file contains the VELO cluster coordinates as well as the Monte Carlo (MC) identifier of the origin particle of this cluster. The file mc_particles.csv associates, for each event, the MC identifier with the properties of the corresponding particle, e.g., momentum, pseudorapidity,, etc.